Two liquid streams are pouring in a container, 8 in. diameter and 15in. high. The mass flow rate of stream A is 1.2 lb/sec and the volume flow rate of B is 0.035 ft3/s. If the specific gravities of A and B are 1.1 and 0.9, respectively, What is the gage and absolute pressure at the bottom of the filled tank in psia.


1
Expert's answer
2021-11-03T06:25:01-0400


In the case of the water stored in a tank, the pressure at its bottom is the weight acting on a unit area of the surface where the tank is kept. To translate that into an equation:

"Pressure = weight/area, and \space weight = mass (m) \times acceleration\space due\space to \space gravity (g)."

This means pressure = m * g/ area.